OpenStudy (arbyj):

PLEASE i need help solving these systems by using substitution. (HELPER WILL GET A METAL) -8x-9y=26 -3x-y=24

OpenStudy (mathmale):

Try this : first, copy down the first equation. Next, multiply the second equation by 9 and write it immediately below the first. Does this result give you an idea of what to do next to solve this system of linear equations by elimination?

OpenStudy (agreene):

substitution means that you put one variable into the other equation. That is: solve one for x or y, and then plug x or y into the other.

OpenStudy (mathmale):

If you are to use substituion, then decide which equation is easier to solve for either x or y, and then solve it for x or y. substitute the result into the other equation.

OpenStudy (arbyj):

yeah, i rearranged the second equation

OpenStudy (arbyj):

do i plug it into y?

OpenStudy (arbyj):

of the first equation?

OpenStudy (anonymous):

\[-8x-9y=26-First~eqn\]\[-3x-y=24-Second~eqn\]\[y=-3x-24-Third~equation\]Use the third equation to substitute into the first equation

OpenStudy (anonymous):

\[-8x-9(-3x-24)=26\]

OpenStudy (anonymous):

\[-8x+27x+216=26\]

OpenStudy (anonymous):

Then subtract both sides by 216

OpenStudy (anonymous):

\[-8x+27x=26-216\]

OpenStudy (anonymous):

\[19x=-190\]Then divide both sides by 19

OpenStudy (arbyj):

wait i thought the 3 became negative?

OpenStudy (anonymous):

\[\frac{ 19x }{ 19 }=-\frac{ 190 }{ 19 }\]\[x=-\frac{ 190 }{ 19 }\]

OpenStudy (arbyj):

scuse me positive

OpenStudy (mathmale):

If you "rearranged the second equation," then y our -3x-y=24 became -3x+3x-y=24+3x, or -y=24+3x. Next, solve this last result for y. I always want to see the actual results you've gotten. Substitute y=-24-3x into the first equation, if y ou haven't already done so. (MARC has done this.) Find x. Then, again using substitution, find y. Finally, check your solutiion (x,y) in both equations to determine whether it satisfies both.

OpenStudy (anonymous):

\[x=-10\]Then substitute x=-10 into the second eqn to solve for y

OpenStudy (anonymous):

\[-3(-10)-y=24\]

OpenStudy (anonymous):

\[30-y=24\]Then subtract 30 on both sides

OpenStudy (anonymous):

\[30-30-y=24-30\]\[-y=-6\]Lastly,divide both sides by -1

OpenStudy (anonymous):

\[\frac{ -1 }{ -1 }y=\frac{ -6 }{ -1 }\]

OpenStudy (anonymous):

\[y=?\]solve it @arbyj

OpenStudy (arbyj):

i got 6

OpenStudy (arihangdu):

Very good your right,

OpenStudy (arbyj):

thank you all for your help :) goodnight

OpenStudy (mathmale):

Please be certain to label each coordinate. Do you mean x=6, or do you mean y=6? Once you have a possible solution, please check it by substitution into the original equations. if both equations come out TRUE, then you have a solution. If not, .... :(

OpenStudy (arbyj):

y=6

OpenStudy (mathmale):

You don't have a solution until you've found both x and y and have checked your x and y into both of the original equations.

OpenStudy (arbyj):

-10,6